OpenStudy (goformit100):

The class mark of the class 90-120 is : (A) 90 (B) 105 (C) 115 (D) 120

OpenStudy (anonymous):

c.m.=upper limit +lowerlimit /2 =90+120/2 =210/2 =105

OpenStudy (goformit100):

Thank you yaar